on a sketch of a mural, 3 inches represent one foot in the mural. A door in the sketch is 2 inches wide by 5 inches high. What is the perimeter of the door in the mural expressed in inches? Show your work.
Answer by mananth(16946)

3 in represents 12 inches
..
2 in will be 2/3 *12 = 8 in
5 in will be 5/3 *12= 20 in
...
perimeter = 2L+2W
Perimeter = 2*8+2*20
Perimeter = 16+40=56 inches
..
m.ananth@hotmail.ca
...